Nov 15, 16:57 UTC
This incident has now been resolved - a postmortem will follow as soon as we have received similar from our operator partner.
Nov 13, 16:47 UTC
We are continuing to monitor for any further issues.
Nov 13, 16:26 UTC
Our monitoring shows the issue has cleared at 16:07 UTC, and services have been stable since then. However, we have not yet received an all clear from the operator partner. We will continue to monitor and update when we have a further update from the operator partner.
Nov 13, 16:25 UTC
Operator partner engineers continue to work on this issue. Further update by 1615 UTC.
Nov 13, 15:45 UTC
Operator partner has advised that engineers are on site and have advised that there is a firewall issue. Next update due from operator partner by 15:30 UTC.
Nov 13, 14:56 UTC
Our operator partner is still working to resolve this issue, and engineers are currently focussing on an issue within their network around VPN termination. They have advised that they will be providing another update within the next 30 minutes.
Nov 13, 14:39 UTC
Our operator partner have advised that they are experiencing an outage and are working to resolve it. More details to follow.
Nov 13, 13:53 UTC
We are investigating an issue with operator partner is currently affecting the following lookup services:
An update will follow shortly.
Nov 13, 13:41 UTC